OpenStudy (anonymous):

when the cost of a car is multiplied by 0.02 the result is 17.50. What is the cost of the car?

OpenStudy (anonymous):

\[x*0.02 = 17.50\] is also equal to \[x*1/5 = 17.50\] just solve for x.

OpenStudy (mathstudent55):

0.02 = 1/50

OpenStudy (anonymous):

whoops ^ mathstudent55 is right, missed a zero